OpenStudy (anonymous):

Find the polynomial function with roots 1, 7, and –3 of multiplicity 2

OpenStudy (anonymous):

(x-1)(x-7)(x+3)^2 =(x^2-8x+7)(x^2+6x+9) =x^4+6x^3+9x^2-8x^3 -48x^2-72x+7x^2+42x+63 =x^4-2x^3-32x^2-30x+63

OpenStudy (anonymous):

te last one is the answer?